How would I make the graphs as well to figure out the empty columns?
Kineties of the Reaction of Phenolphthalein with Nad4 rt Sheet Lab Partner(s): and Caleulations Time (s) Absorbance, A VA .499 120.-1-0. 455 448 150. 180. 5 Run 2 18o. 2 240. C 300. 0 448 240 O S Run 3 360. 480. 6000 S68 720 240. 10.14 Run 4 14-9
media%2Fa22%2Fa22abfa4-1c91-478b-b635-60
Answer #1

The graph of time and absorbance can be simply plotted in microsoft excel.

If an absorbance value is given, then value of ln(A) can be found just by taking the natural logarithm of the absorbance value

For example,

For Run 1, at time t = 0, absorbance, A is 0.999

ln(A) = ln(0.999) = -1.00 x 10-3

Also 1/A = 1/0.999 = 1.001

Run 1 Time (s) Absorbance, A ln(A) 1/A
0. 0.999 -0.0001 1.001
30. 0.985 -0.1511 1.015
60. 0.975 -0.0253 1.026
90. 0.965 -0.0356 1.036
120. 0.955 -0.0460 1.047
150. 0.948 -0.0534 1.055
180. 0.938 -0.0640 1.066